Please answer all parts. No need to explain your answer.

18. (3.46) Here are the IQ test scores of 31 seventh-grade girls in a Midwest school district: 14 100 104 89 102 91 114 114 103 105 108 130 120 132 128 118 119 8672 1 103 74 112 107 103 98 96 112 112 93 Data Set eBook Step 1: We expect IQ scores to be approximately Normal. Make a stemplot to check that there are no major departures from Normality. Use stems with units of tenths. Match your results: 7 24 913 9 68 Stemplot I 10 023334 10 578 11 1122244489 12 08 13 027 24 晑() 8 69 9 13 9 68 Stemplot II 10 23334 10 578 1111222444 11 89 12 12 8 13 27 24 7 8 8 69 9 13 9 68 Stemplot IV 10023334 10 578 1111222444 11 | 89 12 0 12 8 13 02 7 240 8 069 9 1368 Stemplot Ш 101023334578 111122244489 12 08 13 02Step 2: Does your stemplot show departures from Normality? The stemplot shows an uneven distribution with no definite peak. The stemplot is symmetric with two peaks. There are two low outliers, but the stemplot looks reasonably Normal No, the stemplot looks Normal. Step 3: Nonetheless, proportions calculated from a Normal distribution are not always very accurate for small numbers of observations. Calculate the mean for these IQ scores. Give your answer to 2 decimal places. Fill in the blank Step 4: Calculate the standard deviation for these IQ scores. Give your answer to 2 decimal places. Fill in the blank:Step 5: What proportions of the scores are within one standard deviation of the mean? Be careful. This is asking about the data. You will need to use the calculated mean and standard deviation to answer this. Give your answer to 2 decimal places. Fill in the blank: Step 6: What proportions of the scores are within two standard deviations of the mean? Again, this is asking about the data. Give your answer to 2 decimal places. Fill in the blank: Step 7: What would these proportions be in an exactly Normal distribution? 0.68 and 0.997 0.84 and 0.95 0.32 and 0.84 0.68 and 0.95
Answer #1

1)

stemplaot IV

2)

Thus, there are two low outliers, but the stem plot looks reasonably Normal.

3) Mean =Σx/n = 105.84

4) Std. Deviation = √(Σ(x-mean)²/(n-1)) = 14.27

5) within 1 std dev range = (mean-std dev , mean+stdev) = (91.57,120.11)

total number within this range = 23

so, proportion = 23/31 = 0.74

6) within 2 std dev range = (mean-2std dev , mean+2stdev) = (77.2959,134.3815)

total number within this range = 29

so, proportion = 29/31 = 0.94

7) in normal distribution, proportion for within 1 std dev is 0.68 and for 2 std dev is 0.95

so, answer is 0.68 and 0.95